OpenStudy (anonymous):

Can somebody help me with this question? Write an equation in slope-intercept form for a line where slope=0, and that passes through G(3,2)

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

Sub in the given values then solve for y

pooja195 (pooja195):

x1=3 y1=2 m=0

pooja195 (pooja195):

@cobra-strikes are you stuck?

OpenStudy (anonymous):

im plugging in the numbers now lol

OpenStudy (anonymous):

so when plugged in it looks like this right? y-2=0(x-3)

pooja195 (pooja195):

yes :)

pooja195 (pooja195):

now solve for y

OpenStudy (anonymous):

y=2 right?

pooja195 (pooja195):

yes :) good job

OpenStudy (anonymous):

ok so now the equation looks like 2-2=0(x-3) right?

pooja195 (pooja195):

No the final answer is 2...

OpenStudy (anonymous):

so 2=0(x-3)?

pooja195 (pooja195):

The equation itself is y=2

OpenStudy (anonymous):

oh ok thank you!